OpenStudy (wwe123):

the fourth term of GP is greater than the first term,which is positive by 372.The third term is greater than the second by 60.calculate the common ratio and the first term of the progression

OpenStudy (anonymous):

t4 = t1+372 ar^3 = a +372 ...i t3 = t2+60 ar^2= ar +60 ....ii

OpenStudy (anonymous):

If I am not wrong u need to solve that two equations

OpenStudy (wwe123):

but how to solve ?? @sauravshakya

OpenStudy (anonymous):

But I am confused what does by means....

OpenStudy (anonymous):

Did I write the correct equations???

OpenStudy (wwe123):

i am not able to find out a & r in above both equation

OpenStudy (anonymous):

ar^3 = a +372 ar^3 -a =372 a=372 / (r^3 -1) Also, ar^2 =ar +60 ar^2 -ar =60 a= 60 / (r^2 -r)

OpenStudy (anonymous):

So, 372 /(r^3-1) = 60 / (r^2-r) 372/60 = (r^3-1) / (r^2 -r) 372 /60 = (r-1) ( r^2 +r +1) / {r(r-1)} 372/60 = (r^2 +r+1) /r

OpenStudy (anonymous):

Now, can u find r?